OpenStudy (anonymous):

Find the slope and y-intercept of the line. 12x+4y=144

OpenStudy (anonymous):

solve for y 4y=-12x+144 y=-3x+36 slope is -3, y-intercept =36

OpenStudy (anonymous):

y=mx +b where m is the slope and b is the y-intercept

OpenStudy (saifoo.khan):

y-intercept = 144/4 = 36

OpenStudy (anonymous):

y=o, thats how you find the y intercept